Drift Chambers Status Report.

From: Jerzy Cibor (ufcibor@cyf-kr.edu.pl)
Date: Wed Jan 03 2001 - 12:01:14 EST

  • Next message: Flemming Videbaek: "report from Bucharest group"

       The concise status report on drift chambers (as of 12/12/2000).
       ----------------------------------------------
    
    Here is a short description of what we were able to achieve/discover
    working on Brahms drift chambers for a few days in December.
    
    
    ====== T3 ======
    - Two modules were dismounted and three loose wires were fixed. All of
    them were cathode wires. It looked like the wires slipped from the
    solder.
    We attribute these problems to statistics and do hope it won't repeat.
    
    - Having a gas mixture of Ar(60%)+Isobutan(40%) we were able to set 2100
    V
    between anode and cathode. Rapidly increasing current inhibited setting
    higher voltages (in all modules, although the one which was not
    dismounted
    behaved slightly better, which may mean that 36 hours of gas flow is not
    
    enough to establish the equilibrium). A small amount of alcohol added to
    
    the mixture (~2.5%) significantly improved the performance. We were able
    
    to keep 2600 V in all modules with negligible current.
    
    - With a pulse generator we checked all electronics channels, carefully
    investigating those having problems during the last run. We found that
    the
    main sources of the problems were:
     1) bad/loose connections of cables (especially ECL boards and low
      voltage cables)
     2) bad TDC channels
     3) no signals from the chamber (should be further investigated
      whether these channels can be recovered)
    Each problem causes distinctive effects in DC's performance, what helps
    to
    identify it.
    1) and 2) are pretty easy to fix.
    3) regards only less than a dozen channels in the whole T3 (out of ...).
    
    - We looked at cosmics triggering with coincidence in two adjacent CsI
    scintillators placed at the side of T3. From the rough on-line
    estimations
    of the detection efficiency it seems that above 2200 V we entered a
    plateau region. More accurate off-line analysis is in progress.
    
    
    ====== T4 ======
    - All modules keep the voltages (up to 2200 V without alcohol).
    - T4 has been partially connected to multiplexers (we had 8 cards = 256
    channels in hand, the rest is to be shipped to BNL in January).
    - We looked at single channels with a pulser - everything worked as
    anticipated.
    
    
    ====== T5 ======
    - T5 is still waiting for cabling so we were only able to check whether
    it keeps high voltage. Modules 1 and 3 seem to be OK. Module 2 has a
    problem: there is a connection between a cathode wire and a field wire
    in
    one of the planes. Although the module can run as it is, the performance
    
    will suffer and the problem may develop in the future eventually
    disabling
    the module. Therefore it should be dismounted and fixed.
    
    
    
    TO DO list:
    - dismount and fix the middle module of T5
    - check and fix possible connection problems (see description of T3
      status)
    - test different characteristics of DC's (e.g. efficiency vs HV for
      different gas mixtures). Ideally this should be done under real
    working
      conditions i.e. with the beam, but tests with cosmics can provide
      valuable information as well.
    All this will be hopefully done around March.
    
    
    Best Regards,
    Jerzy Cibor and Pawel Staszel
    



    This archive was generated by hypermail 2b29 : Wed Jan 03 2001 - 10:50:03 EST